At Burnt Mesa Pueblo, archaeological studies have used the method of tree-ring dating in an effort to determine when prehistoric people lived in the pueblo. Wood from several excavations gave a mean of (year) 1259 with a standard deviation of 31 years. The distribution of dates was more or less mound-shaped and symmetric about the mean Use the empirical rule to estimate the following. (a) a rangeofyears centered about the mean in which about 60% of the data (tree-ring dates) will be found between and А.D. (b) a rangeof years centered about the mean in which about 95% of the data (tree-ring dates) will be found between and А.D. (c) a range of years centered about the mean in which almost all the data (tree-ring dates) will be found between and A.D
Answer #1

Since the image is not very clear, I am using the below values.

Mean(x) = 1259

Std Dev(s) = 31

a)

Empirical rule states that 68% of the data points lie in the range {x-s, x+s}

= {1259-31, 1259+31} = {1228, 1290}

b)

68% of the data points lie in the range {x-2s, x+2s}

= {1259-2*31, 1259+2*31} = {1197, 1321}

c)

99.7% or ~100% of the data points lie in the range {x-3s, x+3s}

= {1259-3*31, 1259+3*31} = {1166, 1352}